OpenStudy (anonymous):

One thousand raffle tickets are sold for $2.00 each. One grand prize of $500 and two consolation prizes of $100 each will be awarded. Jeremy purchases one ticket. Find his expected value. please show work!

OpenStudy (anonymous):

apparently you add all possibility of prize 500 + 100 +100 =700 there are 1000 tickets 1/1000 *700 =0.7

OpenStudy (anonymous):

ok have a question what about the $2.00 dollars?

OpenStudy (anonymous):

$ 2.00 is what he paid, $.70(our result ) is how much each ticket is worth based on winning chances

OpenStudy (anonymous):

oh okay i see!

OpenStudy (anonymous):

So, probibility wise he would lose 2-0.7 =$1.3

OpenStudy (phi):

http://en.wikipedia.org/wiki/Expected_value The expected value is the winnings weighted by the probability

OpenStudy (phi):

So $500*1/1000 + $200* 2/1000 - $2 *997/1000

OpenStudy (anonymous):

were did u get the 997 from!

OpenStudy (anonymous):

his chance of not winning any thing since 997 tickets don't win anyting

OpenStudy (phi):

3 chances to win something, 997 to lose 3+997=1000

OpenStudy (anonymous):

that thow me off!

OpenStudy (anonymous):

phi are u corrected?

OpenStudy (phi):

??

OpenStudy (anonymous):

what im not if i i actually understand!

OpenStudy (anonymous):

sure

OpenStudy (phi):

Do you have an answer key?

OpenStudy (anonymous):

no

OpenStudy (phi):

Are you studying this or is this a test to see how much math you remember?

OpenStudy (anonymous):

to see how much i remember!

OpenStudy (phi):

Oh oh. There's a lot to remember.

OpenStudy (phi):

Have you ever tried Khan's site? http://www.khanacademy.org/video/expected-value--e-x?playlist=Probability It might help. It's only 15 minutes... just a small bag of popcorn, while you watch.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks phi!